What is the pressure 8000 ft below the surface of the ocean? In SI and English units.

1
Expert's answer
2022-01-10T09:06:56-0500

"p=\rho g h=1000\cdot 9.8\cdot 8000\cdot 0.3048=24~MPa=499~kpsi."